Find the least multiple of 23, which when divided by 18, 21, and 24 leaves the remainder 7,10 and 13 respectively.


1
Expert's answer
2021-10-19T14:39:30-0400

Here

"(18 - 7) = 11, (21 - 10) = 11, (24 - 13) = 11"

LCM of "18, 21" and "24" is "504."

Let the required number be "504k - 11."

Least value of "k" for which "(504k - 11)" is divisible by "23" is "k = 6"

Required number


"= 504 \u00d7 6 - 11\n\n= 3024 - 11\n\n= 3013"